# Quantum enhanced real-time sensing of protein binding process

### v1

@@ III.1 Generating Two-Mode Bright Squeezed State
To generate the Two Mode Bright Squeezed State we use a $\chi^{(3)}$ process called degenerate Four Wave Mixing (FWM) in which annihilation of two photons of same frequency leads to the creation of two new photons of slightly different frequencies. This non-linear process occurs inside the Rb85 atomic vapour contained in a 2.5 cm long cylindrical vapor cell. The atomic ensemble serves as a non-linear medium where the $D_{1}$ line corresponding to the $5S_{1/2}\rightarrow 5P_{1/2}$ transition including the hyper fine structure of the Rb85 atom constitutes a double-$\Lambda$ energy level configuration as shown in Fig 3b.
@@
As shown in the upper half portion of Fig 3a, we use a continuous wave laser of $\lambda=795nm$ which is split into two beams. One of them is used as a strong pump of $450mW$ to drive the FWM process and the other is used as a seed probe which is a cross-polarized weak beam of $150\mu W$ to enhance the FWM process. As explained in Fig 3b, this process requires tuning of both the frequencies where the phase matching condition given by $2\omega_{P}=\omega_{Pr}+\omega_{Cj}$ along with the momentum conservation condition is satisfied the Rubidium cell. The pump frequency $\omega_{p}$ is $\Delta$ detuned from the Doppler broadened excited state which drives the atom from f=2 to the f=3 ground state via stokes scattered photon of frequency $\omega_{Pr}$. Similarly, when another pump photon drives the atom back into the f=2 ground state the anti-stokes scattered photon is generated with frequency $\omega_{Cj}$. This process is enhanced by the usage of the seed probe with the frequency $\omega_{Pr}=\omega_{p}-(\nu_{HF}+\delta)$ where $\nu_{HF}$ is the hyper fine splitting of the electronic ground sate and $\delta$ is the two photon detuning as shown in the energy level diagram.
@@
Single photon detuning $\Delta$ is achieved by locking the main laser by approximately 1 GHZ away from the absorption dip recorded using a reference Rb cell. The Two-photon detuning $\delta$ on the seed probe frequency is achieved by using a double pass AOM of 1.52 GHz. $\delta$ ensures that the time lag between probe and conjugate is very small. The momentum conservation is taken care of by carefully aligning the pump beam and the weak probe beam using appropriate mirror arrangements. Another important parameter necessary for FWM process is the temperature of the cell. The cell window is maintained at around 110 degrees whereas the core temperature is maintained around 90 degrees such that the Rb atoms are contained at the center of the cell to ensure maximum photon atom interaction. This leads to the simultaneously generated, equally polarized probe and the conjugate photons that share quantum correlations and are referred to as being in the Two-Mode Bright Squeezed State.
@@
Thus, all the four control parameters namely the single photon detuning, two photon detuning, the pump, probe alignment, and the temperature of the cell are optimized to achieve the squeezing of 7.8 dB as measured at the source just after the cell. Please refer[35, 36, 33, 37] for more details of generating the TMBSS source using FWM in Rb85 atomic vapour. The flip mirrors in Fig 3a, create a path for the classical beam which technically is the seed probe diverted just before entering the Rb cell. The further arrangements constitute the application of the twin beams in studying the protein binding process.
@@ III.2 Applying TMBSS to the SPR sensor
At the output of the Rb85 cell, we have three beams namely the probe, the conjugate, and the residual pump. As shown in Fig3a, the residual pump is sent to the beam dump using the Glan Taylor beam-splitter which has the extinction ratio of $12\times 10^{5}:$ 1. The twin beams are then aligned and collimated ready to be used along with the SPR sensor to study the protein binding process.
We prepared the sensor using the BK7 (n=1.51) right angle prism from Thorlabs, a microscope glass slide with 50 nm gold layer (SKU: AU.0500.ALSI) bought from Platypus technologies, and the customized flow cell designed and fabricated using on campus 3D printing facility. As shown in Fig 3c, the sample delivery system needs to work such that the proteins get in contact with the gold without leaking from the flow cell. We use O-ring to avoid any leakage from the flow cell as it is tightly clamped against the microscope slide and the prism. The input and the output ports of the flow cell are connected to a peristaltic pump to control the flow of the sample. Another challenge during the clamping process is the formation of air bubbles between the glass slide and the prism. We use immersion oil (56822-50ml Sigma Aldrich, n=1.5) to avoid any air bubbles. We chose the refractive index of the glass of the portion of the microscope slide same as the refractive index of prism as well as the immersion oil such that there is no refraction at the gold dielectric interface.
The sensor is installed on a rotation stage such that the reflectivity can be measured at different incident angles. As shown in the lower half of Fig3a, the probe beam passes through a half wave plate and the PBS so that only the p-polarized light is incident on to the sensor. We place an Acousto-Optic Modulator (AOM) in the probe path, which modulates the intensity to fluctuate at the frequency of 2MHz. This serves as the signal for the experiment. This is an important step to be able to read the signal as well as the noise on the same plot after data processing as explained in the next section. The conjugate beam which is used as a reference, passes through the half-wave plate and the PBS to match the absorption in the probe path. Both the beams are focused on two separate PDA10A2 fixed gain Si detectors from Thorlabs.
@@ III.3 Detection and data processing
Figure 4: (a)Squeezing measurements along the propagation path of the twin beams.(b)The difference in the Standard Quantum Limit and the twin beam noise as measured at point A after post-processing the data using MATLAB
The data was collected using only the oscilloscope and was post processed with MATLAB. Each detector is connected to two channels on the oscilloscope collecting the AC voltage as well as the DC voltage data. We used the DC data mainly to see the absorption dip for water and to visualize the change in absorption during the protein binding process. The AC data which records the fluctuations in the signal as a function of time was converted into frequency domain using the FFT (Fast Fourier Transform) function in MATLAB. This data is used to find the twin beam noise as well as the Shot Noise Level at different points after the Rb cell. The twin beams undergo an absorption of $\alpha_{1}=0.08$ at the window of the cell. The difference between the two noise power as measured right after the cell at point A as shown in 4b was 7.8 dB. Next, the twin beams need to pass through a series of optics before reaching the SPR sensor. Here it looses another $38\%$ of the intensity along with the $20\%$ loss at an off resonance angle inside the sensor giving total absorption $\alpha_{2}=0.5$. The available squeezing at the wings of the absorption dip is around 4.5 dB as measured after the sensor at point B.
Figure 5: AOM modulation to see the signal as well as noise on the same plot
Finally, we measure the squeezing after the sensor at point B as we lock the incident angle where the reflectivity of the prism r is 0.46 $(\alpha=0.54)$ This also constitutes a starting point for the protein biding process. Thus for absorption of $\alpha_{2}=0.71$ we find the squeezing to be around 4dB. Using the data for the noise (both coherent as well as TMBSS) and the AOM modulation as signal we measure the signal to noise ratio as a function of time. As shown in Fig 5 the peak amplitude is taken as the signal and the side bands correspond to the noise for a particular absorption coefficient of the sample.It has been seen before [31] that using the amplitude of the AOM modulation as the signal which is associated to the intensity difference allows us to see the signal and the noise on the same plot plus it helps in removing the technical noise. This forms the complete setup to collect the data involved in studying the biokinetic processes using the Two-Mode Bright Squeezed State along with the Surface Plasmon Resonance sensor.
@@ IV Results
Figure 6: (a)SPR absorption dip (Classical) (b)Protein binding curve (Classical)
@@
Next, we lock the angle at $\phi=32$ degrees which is 0.8 degrees left of the resonance angle $\theta_{r}$ corresponding to absorption $\alpha=0.59$ absorption or the reflection coefficient $r=0.41$. We then filled the cavity with $10\%$ BSA solution and recorded the change in reflectivity as the protein binds the gold film while changing the refractive index of the sample. As mentioned earlier the angle of resonance shifts towards the right as the refractive index increases, which happens because of the protein binding process. Thus, fixing the angle on the left side of the dip results in an increase of the reflected intensity as the protein binds. This could be visualized in Fig 6b, where the reflected intensity slowly increases with time and the slope of the curve gives the rate of binding. The error bars in the plot correspond to 10 consecutive readings taken at the same time stamp.
Figure 7: (a)Squeezing as a function of incident angle following the SPR dip.(b)Squeezing as a function of reflectivity.
Another important parameter helpful in deciding the angle is squeezing. Squeezing in dB as defined by Eq. (6) depends on the absorption that changes at different incident angles on the SPR sensor as shown in Fig7. We experimentally collected the Noise power at different angles around the absorption dip and compared it with the SNL at each point. The SNL was recorded for one angle and using the linear relation of the noise of the coherent source $(\Delta N)^{2}\propto\left\langle N\right\rangle$ we calculated the SNL for the rest of the points. As shown in Fig 7a, we get squeezing of around 4 dB at an angle of 0.8 degrees left of $\theta_{r}$ as decided using the classical SPR dip data. The squeezing as a function of reflectivity is reported in Fig7b.
Figure 8: (a)BSA binding with TMBSS (Quantum) (b)SNR comparison between classical and quantum
Finally, we plot the data for the protein binding process using the refelctivity of the probe beam of the TMBSS source. This plot is shown in fig 8a. As the protein binds, the signal increases due to the shift in SPR dip towards the right increasing the reflectivity similar to the coherent source. While taking the data, we change the conjugate manually by the same proportion as the probe. It more or less ensures that we have the same absorption on conjugate as the probe. Thus we ensure that we get the maximum quantum advantage possible at that absorption level.
It is easier to see the quantum advantage in the SNR plot as a function of time reported in fig 8b. The quantum advantage here is defined as,
@@
The quantum advantage of the TMBSS light is realized when we take the intensity difference between the probe and the conjugate. The signal in this plot is the amplitude of the modulated peak at 2MHz and the noise is quantified by taking the average of the side bands as shown in Fig 5. We can see the quantum advantage of 4dB despite the total absoeption of $\alpha=0.74$. The SNR for TMBSS looks flat because of a very small change in absorption. This plot demonstrates the quantum advantage of using TMBSS in applications involving biokinetic processes.
@@ V Conclusion
In summary, we have demonstrated the quantum advantage of using Two Mode Bright Squeezed State (TMBSS) in improving the sensitivity of a Surface Plasmon Resonance (SPR) based sensor to study biokinetic processes. We used the protein – Bovine Serum Albumin (BSA) binding to the gold film on the sensor as an example of such biokinetic process. As the protein binds, the reflectivity of the sensor goes up due to the shift in surface plasmon resonance angle. We showed 4dB of squeezing (difference between the coherent source noise and the twin beam noise) as we recorded the signal-to-noise ratio as the function of time and is maintained throughout the binding process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of $74\%$ from the source until the final detection after the sensor.
Our work reinforces the utility of TMBSS source highlighting its primary advantage: the continuous variable nature as opposed to the single photon sources that allows for a greater number of photons while still employing the quantum mechanical benefits that arise from the intensity difference squeezing. TMBSS constitutes a readily available quantum probe which is better than the coherent source because of its sub-shot noise characteristic thus allowing for lower intensities with comparable results, making it more gentle to the biological samples. One of the of the main challenges we encountered in the implementation of this technology was the stabilization issue with the twin beams. Thus, if the source is stabilized and when the both the beams are controlled in such a way that the ratio between probe and conjugate stays the same, this study can be extended to measure the rate of protein binding with higher precision than current technique involving a coherent source. It can also be used to study protein-protein, protein-ligand biding processes that are very significant in the field of biochemistry and biophysics.
Overall, this project brings together two state-of-the-art techniques and provides a most practical, quantum-enhanced SPR setup for improving the sensitivity of time-dependent measurements in the study of biological processes.

### v2

@@ III.1 Generating Two-Mode Bright Squeezed State
To generate the Two Mode Bright Squeezed State we use a $\chi^{(3)}$ process called degenerate Four Wave Mixing (FWM) in which annihilation of two photons of same frequency leads to the creation of two new photons of slightly different frequencies. This non-linear process occurs inside the Rb85 atomic vapor contained in a 2.5 cm long cylindrical vapor cell. The atomic ensemble serves as a non-linear medium where the $D_{1}$ line corresponding to the $5S_{1/2}\rightarrow 5P_{1/2}$ transition including the hyperfine structure of the Rb85 atom constitutes a double-$\Lambda$ energy level configuration as shown in Fig. 3(b).
@@
As shown in the upper half portion of Fig. 3(a), we use a continuous wave laser of $\lambda=795nm$ which is split into two beams. One of them is used as a strong pump of $450mW$ to drive the FWM process and the other is used as a seed probe which is a cross-polarized weak beam of $150\mu W$ to enhance the FWM process. As explained in Fig. 3(b), this process requires tuning of both the frequencies where the phase matching condition given by $2\omega_{P}=\omega_{Pr}+\omega_{Cj}$ along with the momentum conservation condition is satisfied the Rubidium cell. The pump frequency $\omega_{p}$ is $\Delta$ detuned from the Doppler broadened excited state which drives the atom from f=2 to the f=3 ground state via stokes scattered photon of frequency $\omega_{Pr}$. Similarly, when another pump photon drives the atom back into the f=2 ground state the anti-stokes scattered photon is generated with frequency $\omega_{Cj}$. This process is enhanced by the usage of the seed probe with the frequency $\omega_{Pr}=\omega_{p}-(\nu_{HF}+\delta)$ where $\nu_{HF}$ is the hyperfine splitting of the electronic ground sate and $\delta$ is the two photon detuning as shown in the energy level diagram.
@@
Single photon detuning $\Delta$ is achieved by locking the main laser by approximately 1 GHz away from the absorption dip recorded using a reference Rb cell. The Two-photon detuning $\delta$ on the seed probe frequency is achieved by using a double pass AOM of 1.52 GHz. $\delta$ ensures that the time lag between probe and conjugate is very small. The momentum conservation is taken care of by carefully aligning the pump beam and the weak probe beam using appropriate mirror arrangements. Another important parameter necessary for FWM process is the temperature of the cell. The cell window is maintained at around 140°C to prevent the Rb atoms from settling on the window whereas the core temperature is maintained around 90°C as required for the process of FWM. This leads to the simultaneously generated, equally polarized probe and the conjugate photons that share quantum correlations and are referred to as being in the Two-Mode Bright Squeezed State.
@@
Thus, all the four control parameters namely the single photon detuning, two photon detuning, the pump, probe alignment, and the temperature of the cell are optimized to achieve the squeezing of 7.8 dB as measured at the source just after the cell. Please refer[38, 39, 36, 40] for more details of generating the TMBSS source using FWM in Rb85 atomic vapor. The flip mirrors in Fig. 3(a), create a path for the classical beam which technically is the seed probe diverted just before entering the Rb cell. The further arrangements constitute the application of the twin beams in studying the protein-gold binding process.
@@ III.2 Applying TMBSS to the SPR sensor
At the output of the Rb85 cell, we have three beams namely the probe, the conjugate, and the residual pump. As shown in Fig. 3(a), the residual pump is sent to the beam dump using the Glan Taylor beam-splitter which has the extinction ratio of $12\times 10^{5}:$ 1. The twin beams are then aligned and collimated to be used along with the SPR sensor to study the protein binding process.
As shown in the lower half of Fig. 3(a), the probe beam passes through a half wave plate and the Polarizing Beam Splitter (PBS) so that only the p-polarized light is incident on the sensor. We place an Acousto-Optic Modulator (AOM) in the probe path, which modulates the intensity to fluctuate at the frequency of 2MHz. This is an important step to be able to obtain the signal as well as the noise on the same plot after data processing as explained in the detection and data processing section.
Next, the probe beam reaches the sensor used to study the BSA protein adsorption on gold film. The sensor is assembled using commercially available SPR chip along with the on-campus facilities as explained in appendix B. The flow cell cavity as shown in Fig. 3(c) is filled up with the required sample using the tubes connected via a peristaltic pump. In this paper, we use $10\%$ BSA solution which was prepared by dissolving 1gm Bovine Serum Albumin (BSA) Lyophilized powder into 10 ml of deionized water. We wait for approximately one hour to ensure complete dissolution and uniformity of the solution before injecting the fluid into the flow cell. The sensor assembly is then installed on a rotation stage such that the reflectivity can be measured at different incident angles. All the angles are measured using the incident angle of $\theta=45\degree$ as the reference angle which is experimentally fixed using two irises.
The conjugate beam which is used as a reference, passes through the half-wave plate and the PBS to match the absorption in the probe path. Both the beams are focused on two separate PDA10A2 fixed gain Si detectors from Thorlabs.
@@ III.3 Detection and data processing
Figure 4: (a) Squeezing measurements along the propagation path of the twin beams. (b) The difference in the Standard Quantum Limit and the twin beam noise as measured at point A after post-processing the data using MATLAB.
The data was collected using only the oscilloscope and was post processed with MATLAB. Each detector is connected to two channels on the oscilloscope collecting the AC voltage as well as the DC voltage data. We used the DC data mainly to see the absorption dip for water and to visualize the change in absorption during the protein binding process. The AC data which records the fluctuations in the individual signals as a function of time was converted into frequency domain using the FFT (Fast Fourier Transform) function in MATLAB. This data is used to find the twin beam noise as well as the Shot Noise Level at different points after the Rb cell. The twin beams undergo an absorption of $\alpha_{1}=0.08$ at the window of the cell. The difference between the two noise power as measured right after the cell at point A as shown in Fig. 4(b) was 7.8 dB. Next, the twin beams need to pass through a series of optics before reaching the SPR sensor. Here it looses another $38\%$ of the intensity along with the $20\%$ loss at an off resonance angle inside the sensor giving total absorption $\alpha_{2}=0.5$. The available squeezing at the wings of the absorption dip is around 4.5 dB as measured after the sensor at point B.
Figure 5: AOM modulation to view the sensor signal as well as noise on the same plot.
Finally, we measure the squeezing after the sensor at point B as we lock the incident angle where the reflectivity of the prism r is 0.46 $(\alpha=0.54)$. This also constitutes a starting point for the protein biding process. Thus for absorption of $\alpha_{2}=0.71$ we find the squeezing to be around 4 dB. Using the data for the noise (both coherent as well as TMBSS) and the AOM modulation as sensor signal we measure the signal to noise ratio as a function of time. As shown in Fig. 5 the peak amplitude is taken as the signal and the side bands correspond to the noise for a particular absorption coefficient of the sample. It has been shown before [31] that using the amplitude of the AOM modulation as the signal which is associated to the intensity difference allows us to view the signal and the noise on the same plot plus it helps in removing the technical noise. This forms the complete setup to collect the data involved in studying the kinetics of the biological processes using the Two-Mode Bright Squeezed State along with the Surface Plasmon Resonance sensor.
@@ IV Results
Figure 6: (a) The reflectivity of the sensor is recorded using the classical beam in the probe path. The scanning angle plotted on x-axis is measured with respect to the angle of resonance for deionized water such that zero corresponds to the incident angle $\theta=66.02\degree$. (b) The reflectivity of the sensor is recorded against time keeping the scanning angle fixed at 0.8 degrees left of the resonance angle. The reflected intensity increases as the BSA protein binds the gold surface on the sensor.
@@
Next, we lock the scanning angle at $32\degree$ to the normal at the air-prism interface (0.8 degrees left of the resonance angle) corresponding to absorption $\alpha=0.59$ absorption or the reflection coefficient $r=0.41$. We then fill the flow cell cavity with $10\%$ BSA solution that was prepared as explained in the experimental setup section. We then record the change in reflectivity as the protein binds the gold film while changing the local refractive index of the sample. As mentioned earlier the angle of resonance shifts towards the right as the refractive index increases resulting from the protein-gold binding process. Fixing the angle on the left side of the dip leads to an increase of the reflected intensity as the protein binds. This could be visualized in Fig. 6(b), where the reflected intensity slowly increases with time and the slope of the curve gives the rate of binding. The error bars in the plot correspond to 10 consecutive readings taken at the same time stamp.
Figure 7: (a) Squeezing was measured at different scanning angles near the absorption dip. (b) This plot shows the variation of squeezing with reflectivity, providing a clearer visualization of the expected squeezing as BSA binds to the gold. The data was derived using the reflectivity versus angle plot.
Squeezed light characterization: Squeezing in dB as defined by Eq. (5) depends on the absorption that changes at different incident angles on the SPR sensor as shown in Fig.7(a). We experimentally collected the noise power at different angles around the absorption dip and compared it with the SNL at each point. The SNL was recorded for one angle and the rest of the points were calculated using the Poisson characteristic of the coherent source $(\Delta N)^{2}=\left\langle N\right\rangle$. The coherent source is verified to be shot-noise limited around the working frequency of 2MHz as shown in Appendix C. As shown in Fig. 7(a), we get squeezing of around 4 dB at an angle of 0.8 degrees left of $\theta_{r}$ as decided using the classical SPR dip data. The squeezing as a function of reflectivity is reported in Fig.7(b).
Figure 8: (a) The reflectivity of the sensor is recorded using the TMBSS beam in the probe path. The reflected intensity increases as the BSA protein binds the gold surface on the sensor. (b) The signal-to-noise ratio calculated as explained in the detection and data processing section is plotted against time as the BSA protein binds to the gold. 4dB of quantum advantage is maintained through out the process.
Quantum results: We plot the data for the protein binding process using the reflectivity of the probe beam of the TMBSS source. This plot is shown in Fig. 8(a). As the protein binds, the signal increases due to the shift in SPR dip towards the right increasing the reflectivity similar to the coherent source. While taking the data, we change the conjugate manually by the same proportion as the probe. Matching the absorption on conjugate path with that of the probe path ensures that we get the maximum quantum advantage possible at that absorption level.
It is easier to see the quantum advantage in the SNR plot as a function of time reported in Fig. 8(b). The quantum advantage here is defined as,
@@
The quantum advantage of the TMBSS light is realized when we take the intensity difference between the probe and the conjugate. The signal in this plot is the amplitude of the modulated peak at 2 MHz and the noise is quantified by taking the average of the side bands as shown in Fig. 5. We can see the quantum advantage of 4 dB despite the total absorption of $\alpha=0.74$. The SNR for TMBSS looks flat because of a very small change in absorption. This plot demonstrates the quantum advantage of using TMBSS in applications involving biokinetic processes.
Figure 9: The nonlinear fit function was used on the probe data to extract the mean value of the parameter $k_{s}$.
Kinetic parameter estimation sensitivity: The interaction dynamics of the adsorption of BSA protein molecule on the gold surface and the corresponding reflectivity change can be modeled based on the framework given in [41]. The Reflected intensity denoted by R can be expressed as a function of t as follows,
$R(t)=A+B\left(1-e^{-k_{s}t}\right)$ (7)
where $A$ is the initial reflected intensity at time $t=0$, $B$ is the maximum change in reflected intensity as it reaches equilibrium and $k_{s}$ is the observable rate constant in units of $s^{-1}$. The approximate mean value of the rate constant $k_{s}$ can be found out using the non-liner fit as shown in Fig. 9. For our setup, we find the values as follows $A=0.4601$, $B=0.0603$ and $k_{s}=0.0046\,s^{-1}$. Here we focus mainly on showing the improvement in the sensitivity of estimating the parameter $\Delta k_{s}$ when the signal is shot-noise limited. Eq. (8) can be written as,
$\frac{R-A}{B}=1-e^{-k_{s}t}$ (8)
Thus the uncertainty in R for the coherent case is given by,
$\frac{(\Delta R)_{Coh}}{B_{Coh}}=e^{-k_{s}t}(\Delta k_{s}t)_{Coh}$ (9)
Using Eq.(9) we can write,
$(\Delta R)_{Coh}=-\left(R+A-B\right)_{Coh}(\Delta k_{s}t)_{Coh}$ (10)
We can write the same equation for the TMBSS case. Knowing that A,B are constants and that we are comparing the quantum and classical cases for same intensities R for a given time instant we can take the ratio to get,
$\frac{(\Delta R)_{Coh}}{(\Delta R)_{TMBSS}}=\frac{(\Delta k_{s})_{Coh}}{(\Delta k_{s})_{TMBSS}}$ (11)
The noise reduction in TMBSS case is attributed to the intensity difference squeezing between the probe and the conjugate whereas the mean value of signal is same for both cases. Hence the ratio in Eq.(12) could be written in terms of the Signal-to-Noise Ratio (SNR) as,
$\frac{(\Delta k_{s})_{Coh}}{(\Delta k_{s})_{TMBSS}}=\frac{(SNR)_{TMBSS}}{(SNR)_{Coh}}$ (12)
Using the quantum result as shown in Fig. 8(b) the 4dB improvement in SNR implies the sensitivity of the estimation of $k_{s}$ is improved by $60\%$.
@@ V Conclusion
In summary, we have demonstrated the quantum advantage of using Two Mode Bright Squeezed State (TMBSS) in improving the sensitivity of a Surface Plasmon Resonance (SPR) based sensor to study biokinetic processes. As an example of one such process, we used the protein – Bovine Serum Albumin (BSA) binding to the gold film on the surface of the sensor. As the protein binds, the reflectivity of the sensor goes up due to the shift in surface plasmon resonance angle. We showed 4dB of squeezing (difference between the coherent source noise and the twin beam noise) as we recorded the signal-to-noise ratio as the function of time and is maintained throughout the binding process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of $74\%$ from the source until the final detection after the sensor.
We successfully demonstrated the reduction in the state noise of the source which contributes to a $60\%$ improvement in the sensitivity of estimation of the binding parameter $k_{s}$.Thus, integration of a well stabilized source of TMBSS with SPR can achieve more precise measurements of the rate of protein binding to gold surpassing the limitations of the current shot-noise limited techniques. It can also be used to study protein-protein, protein-ligand biding processes that are very significant in the field of biochemistry and biophysics.
Our work reinforces the utility of TMBSS source highlighting its primary advantage: the continuous variable nature as opposed to the single photon sources that allows for a greater number of photons while still employing the quantum mechanical benefits that arise from the intensity difference squeezing. TMBSS constitutes a readily available quantum probe which is better than the coherent source because of its sub-shot noise characteristic thus allowing for lower intensities with comparable results, making it more gentle to the biological samples. The combination of TMBSS along with the sub-diffraction-limited sensing property of the SPR presents a highly promising approach in the realm of cellular biology.
Overall, this project brings together two state-of-the-art techniques and provides a most practical, quantum-enhanced SPR setup aimed at improving the sensitivity of time-dependent measurements in the study of biological processes.